MICQ(1) MICQ(1) NAME mICQ−asmall,textmodeICQclient SYNOPSIS micq[-h][-c][-b<basedir>][-i<locale>][-v[level]][-l<logplace>][-u<UIN>][-p<passwd>][-s <status>][-C<commands>] DESCRIPTION mICQisatextbasedICQclientsupportingmanyoftheICQfeaturesprovidedbytheclientby Mirabilis.mICQallowstochangetheuserpassword,createnewaccounts,updatinguserinfo, searchotherusers,sendSMS,opendirectconnections,sendacknowledgedmessages,transferfiles, detectotheruser’sclient,sendUTF-8encodedmessages,executearbitrarycommandsuponany event,etc. OPTIONS −h|−−help Displayashorthelptextandexit. −c|−−nocolor Disabletheuseofcolor. −b|−−basedir<basedir> SpecifybasedirasthebasedirectorywheremICQwilllookforitsfiles,inparticularits micqrcfile. −i|−−i18n<locale> Usethelocalelocaleinsteadoftheonederivedfromenvironmentvariables. −v|−−verbose[level] Setverbosemode,orsetverbositytolevel. −l|−−logplace<file>|<dir> Usefileasfiletowriteloginto,orusedirasadirectorytocreateperuserlogfilesin. −u|−−uin<uin> Loginwithuinonly.Mayberepeated. −p|−−passwd<passwd> Loginwithpasswdaspassword.Mayberepeatedforeach-uswitch,orsingle. −s|−−status<status> Loginwithstatusasstatus,whichmaybegivennumericallyorinv,dnd,occ,na,away, ffc,oroff[line]. −C|−−cmd<commands> ExecutemICQcommandscmdsafterlogininsteadofthedefaultegcommand. USAGE Uponfirststart(forthegivenbasedirectory),mICQwillaskforaUIN,orcreateanewone. Afterthefirstlogin,theserverbasedcontactlistwillbeimportedautomatically. Basiccommandsincludemsgtomessageothercontacts(trypressingthe<tab>keyseveraltimes onanemptyinputlinetobringupmsgcommandstomessagecontactsrecentlytalkedto),addto addnewcontacts,update,other,andabouttochangeyourmetadataontheICQserver,and,of course,helpforonlinehelp.Acompletelistofcommandsisavailableinthemicq(7)manpage. FEATURES mICQisveryportableandrunsonLinux,DOS,OS/2,Win32,BeOS,AmigaOS,GNUHURD andanyothermodernunixsystem.mICQversion0.5isknowntorunonLinux,AIX,HPUX, Tru64,AmigaOS,partiallyBeOSandotherunices.However,theotherportsmayormaynotstill workandanyonewishingtotestthemiswelcometo. HISTORY mICQ 1 MICQ(1) MICQ(1) 16Aug1998 MatthewD.Smithreleasesmicquntotheworld.Followedby12datedreleasesuntil25 Sept1998. 05Jan2001 MatthewD.Smithdiedinacaraccident.UserspickupwhereMatthewleftoff,formalizemICQundertheBSDlicenseandrestartthedevelopmentofmICQ. 10June2002 Version0.4.9isreleasedundertheGPLlicence,finallyimplementingthenewversion8of theICQprotocol,donebyRu¨ digerKuhlmann. MAILING-LIST ThemICQmailinglistismicq-list@micq.org.Informationaboutsubscribingandarchivesisavail-ableathttp://www.micq.org/pipermail/micq-list/. DOWNLOAD ViaHTTPfromhttp://www.micq.org/download. TogetthenewestversionfromCVS,enter: cvs−d:pserver:anonymous@micq.org:/var/lib/cvscomicq SEEALSO micqrc(5),micq(7) FILES ˜ /.micqrc OldconfigurationfileformICQ−obsolete. ˜ /.micq/ ThedefaultbasedirectorywheremICQlooksforitsfiles. BASE/micqrc ConfigurationfileformICQ.Containsglobalconfiguration. BASE/status ConfigurationfileformICQ.Containscontactslistandcontactoptions. BASE/i18n/*.i18n LocaloverridesformICQ’stranslationfiles. ˜ /.micq/i18n/*.i18n LocaloverridesformICQ’stranslationfiles. BASE/contacts/ Directorywheremetadataforcontactsisstored. BASE/history/ Defaultdirectoryforlogfiles. BASE/files/ Directorywhereincomingfiletransfersaresaved. BASE/remote-control OlddefaultlocationoftheFIFOusedforscripting.Renamedasitconfusedlusers. BASE/scripting DefaultlocationoftheFIFOusedforscripting. PREFIX/share/micq/*.i18n InternationalizationfilescontainingtranslationsofallstringsmICQoutputs. BUGS SendbugreportsorcommentstoRu¨ digerKuhlmann<micq@ruediger-kuhlmann.de>(inGerman orEnglish). mICQ 2 MICQ(1) MICQ(1) AUTHOR ThismanualpagewasderivedbyJamesMorrison<ja2morrison@student.math.uwaterloo.ca>from theminimalmanpagewrittenbyJordiMallach<jordi@sindominio.net>fortheDebian GNU/Linuxsystemandlaterkeptup-to-dateby<micq@ruediger-kuhlmann.de>. mICQ 3